OpenStudy (aaronandyson):

sec^2 theta = 2 + 2 tan theta,find tan theta

OpenStudy (aaronandyson):

@MrNood

OpenStudy (owlcoffee):

so, an identity proof.

OpenStudy (aaronandyson):

I GOT sec^4 theta - 4sec theta - 4 sec^2 theta + 4 = 0

OpenStudy (owlcoffee):

sre you sure you typed the exercise correctly?

OpenStudy (aaronandyson):

The question?Yes.

OpenStudy (owlcoffee):

I believe there should be something else in that one.

OpenStudy (aaronandyson):

\[\sec^2 \theta = 2 +2\tan \theta\]

OpenStudy (owlcoffee):

I end up getting 2 sec^2 theta

OpenStudy (aaronandyson):

What mistake are we making?

OpenStudy (aaronandyson):

Can I please see your work?

OpenStudy (owlcoffee):

I thin I did a wrong thing.

jhonyy9 (jhonyy9):

so we know that sec theta = 1/cos theta than sec^2 theta = 1 / cos^2 theta = (sin^2 theta + cos^2 theta )/cos^2 theta = =sin^2 theta /cos^2 theta +cos^2 theta /cos^2 theta = tan^2 theta +1 continuing this way i think we can getting this needed 2+2tan theta @ganeshie8 ?

OpenStudy (owlcoffee):

well tan^2 theta +1 is already sec^2 theta

ganeshie8 (ganeshie8):

tan^2 theta +1 = 2+2tan theta Nice! I think that is just a quadratic which can be solved easily !

OpenStudy (owlcoffee):

Isn't it an identity proof?

OpenStudy (sachintha):

I too thought it was an identity proof. Are you looking for general solutions?

OpenStudy (mayankdevnani):

@Owlcoffee it isn't He forgot to mention,find tan theta

jhonyy9 (jhonyy9):

yes @ganeshie8 i think you know my knowledge better from all these OS users - and thank you your opinion - have a nice day ! to everybody

OpenStudy (owlcoffee):

ah, I misread. Then starting from: \[\tan^2 \theta +1 =2+2\tan \theta \iff \tan^2 \theta -2\tan \theta -1=0\] calling tan theta = g \[g^2-2g-1=0 \] \[g_1=1-\sqrt 2 \] \[g_2 = 1+ \sqrt 2 \]

OpenStudy (aaronandyson):

How did I get 1 and -1 then?

OpenStudy (aaronandyson):

@Owlcoffee

OpenStudy (owlcoffee):

just moving everything to the LHS

OpenStudy (aaronandyson):

I also got g^2 - 2g - 1 = 0

OpenStudy (owlcoffee):

Yes, now you have to solve it as you would solve any quadratic equation.

OpenStudy (aaronandyson):

I got 1 and -1

OpenStudy (owlcoffee):

I put the solutions for it above, the results are, if you apply the quadratic formula: \[g=\frac{ 2 \pm \sqrt{2^2-4(1)(-1)} }{ 2 }\] \[g=\frac{ 2 \pm \sqrt{8} }{ 2 } \iff g=\frac{ 2 }{ 2 } \pm \frac{ 2\sqrt2 }{ 2 }\] \[g=1\pm \sqrt 2\]

OpenStudy (owlcoffee):

Just deconstruct the change of variables and you'll obtain the result.